OpenStudy (anonymous):

A person is considered to have a certain symptom when his or her temperature is higher than 100.94 degrees Fahrenheit. The formula F=9/5C+32 can be used to convert Celsius temperatures C to Fahrenheit temperatures F. For which Celsius temperatures is the person considered to have the symptom?

OpenStudy (anonymous):

Thnk u Ham, but I knew that.. .. I have to show my work, and I keep going wrong somewhere. .. =(

OpenStudy (hamsandwich):

hmm... not too sure how to do this without a converter. I can try though. First, multiply everything by 5. You get 5f = 9c + 160. Subtract 160 from both sides and you get 5f - 160 = 9c.

OpenStudy (hamsandwich):

Since you know what f is (100.94), plug that in for f. You then get 507.4 - 160 = 9c.

OpenStudy (hamsandwich):

507.4 - 160 = 347.4. So, 347.4 = 9c. Then, divide everything by 9, and you get 38.6. Apparently google was off by .3 degrees.

OpenStudy (anonymous):

Thank you!! Thank you so very much!! I really appreciate your help!! <3

OpenStudy (hamsandwich):

you're welcome ♥